### v1

# Quasicrystals as an intermediate form of matter between crystalline and amorphous solids

Abstract
Quasicrystals have been observed in a variety of materials ranging from metal alloys to block copolymers and represent an “intermediate” form of matter between crystals and amorphous materials (glasses and liquids) in that their structural and dynamical properties can not readily described in terms of conventional solid-state models of liquids and solids. In the present work, we present a comprehensive analysis of basic thermodynamic and dynamic properties of quasicrystals to better understand the nature of the atomic motion underlying diffusion and structural relaxation in these materials. As our model system, we investigate a dodecagonal quasicrystal using molecular dynamics (MD) simulations in two dimensions (2D), subject to periodic boundary conditions. We observe a two-stage relaxation dynamics in the self-intermediate scattering function $F_{s}(k,t)$ of our quasicrystal material involving a fast $\beta$-relaxation on a ps timescale and $\alpha$ relaxation process having a highly temperature dependent relaxation time whose activation energy varies in concert with the extent of string-like collective motion. Multi-step relaxation of the intermediate scattering function and string-like collective atomic motion have similarly been observed ubiquitously in glass-forming liquids at low temperatures and in crystalline materials at elevated temperatures where structural relaxation and diffusion are both non-Arrhenius. After examining the dynamics of our quasi-crystalline material in great detail, we conclude that its dynamics more closely resemble observations on metallic glass-forming liquids, in qualitative accord with previous neutron scattering studies.
I Introduction
According to the traditional paradigm of solid-state physics, solids are divided into two types: crystals and amorphous (or non-crystalline) materials. Crystals display long-range order and they are defined by an ordered lattice with translational periodicity. They are compatible only with discrete $1$-, $2$-, $3$-, $4$-, and $6$-fold rotational symmetries and they present sharp Bragg peaks in x-ray diffraction. On the contrary, amorphous solids lack all of these properties, and in particular, they do not exhibit long-range order. In April 1982, Shechtman [1] (see also Ref. 2) discovered a crystal exhibiting $5$-fold symmetry – a type of ordered structure previously considered to be “impossible”. Despite lacking periodicity, the icosahedral phase with $5$-fold symmetry was shown to exhibit well-defined and bright Bragg peaks as for crystals, and differently from amorphous solids. The crystal that Shechtman discovered was the first example of an aperiodic crystal, or as latter termed by Levine and Steinhardt [3, 4] a “quasicrystal”, short for quasi-periodic crystal [5, 6, 7, 8]. It was recognized early by Ishimasa et al. [9] that quasicrystals represent an intermediate ordered state between crystalline and amorphous matter (see Fig. 1), which is not periodic, i.e., quasi-periodic.
Figure 1: Quasicrystals as an intermediate form of matter.
Angell [10] has argued that many forms of other matter likewise qualify as being called intermediate forms of matter, e.g., proteins, plastic crystals, incommensurate structures, twisted 2D heterostructures and glass-formers undergoing fragile-to-strong glass formation, and he observed that such materials often seemed to share common thermodynamic and dynamic characteristics, even if the fundamental origin of these properties is not currently understood theoretically. In the case of quasi-crystals, neutron scattering studies have repeatedly indicated that the properties of quasi-crystalline materials and incommensurate structures often resemble those of metallic glasses rather than crystalline materials. This was first noticed for the static structure factor [11], and then recognized also for dynamical properties, connected to anomalies in the vibrational density of states (vDOS) and the low-temperature heat capacity [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. Given the relatively ordered nature of quasi-crystalline materials in relation to metallic glass materials, it is then interesting to take a closer look at the dynamics of these materials and to compare with established trends for metallic glasses and heated crystalline materials to determine if this similarity to glass-forming materials also holds in relation to structural relaxation and atomic diffusion. We next review some essential properties of quasi-crystalline materials.
Mathematically, every aperiodic structure in $d$ dimensions can be obtained using projection/cut operators from a periodic structure in a higher-dimensional space (the “superspace”), which is equivalent to saying that a basis of $d$ independent unit vectors $\hat{x}_{n}$ is incomplete to describe the reciprocal lattice in a quasicrystal. The number of additional fundamental vectors needed to determine the position of the lattice points in an aperiodic structure coincides exactly with that of the extra-dimensions in the superspace [25, 26]. This additional extra-dimensional freedom has a theoreticallly predicted consequence. In quasicrystals, one expects low-energy excitations additional to phonons, termed “phasons” [27]. These predicted excitations are associated with infinitesimal rigid translations along these extra directions [28, 29, 30, 31, 32, 33, 34, 35]. But their stability and observability are outstanding questions. From hydrodynamics and effective field theory, one can formally derive that at long-wavelengths phasons are not propagating waves but rather overdamped diffusive modes [36, 37, 38, 39, 34, 35]. This situation changes for large wave-vectors, where the propagating character is restored following a telegrapher-type dynamics [40]. Both these regimes have been claimed to be observed [41, 42, 43], especially in incommensurate structures [44, 45]. The possible diffusive nature of the phason is related to the observation that rigid translations in the extra-dimensional superspace do not correspond to vibrational displacements in real space, but rather to atomic re-arrangements and particle jumps. Despite these widespread concepts, some authors [46] (see also Ref. 47 for a discussion) have questioned the microscopic origin of the phason modes and have indicated that there is no physical evidence for the existence of long wavelength phason fluctuations, despite the observation of this collective atomic motion.
On the other hand, the problem of bulk self-diffusion in quasicrystals has attracted a lot of attention in the past as a possible smoking gun for the existence of phasons. In 1993, Kalugin and Katz [48] proposed a novel mechanism for diffusion in quasicrystals based on phason dynamics. More concretely, they argued that in quasicrystals the presence of vacancies is not necessary for atomic jumps since the latter are naturally provided by the phason degrees of freedom. The temperature ($T$) dependence of the diffusivity in quasicrystals should then exhibit a deviation from a straight line in the Arrhenius plot, which is the case for vacancy-mediated diffusion in ordered crystals. A first realization of this proposal was observed by Joseph et al. [49] using Monte Carlo simulations. A two-phase diffusive dynamics, interpolating between different thermally activated regimes, was experimentally observed in Ref. 50 using radiotracer techniques (see also Ref. 51), where it was proposed that high-temperature diffusion in quasicrystals is dominated by vacancies while the low-temperature one by phason modes. The same observations were reported using NMR (Nuclear Magnetic Resonance Spectroscopy) experiments [52, 53] and quasi-elastic neutron scattering data [54] (see also Ref. 55). Phason defects have been also identified using scanning tunneling microscopy and their dynamics is related to the static stress field of dislocations [56]. Despite all these observations, it was pointed out by Dolinsek et al. [57] (see also Comment by Coddens [58]) that, although diffusive atomic jumps in quasicrystals exhibit properties compatible with the elementary excitations in the form of phason jumps, very similar atomic motion is observed also in “vacancy-ordered” bcc (body center cubic) crystals. This poses serious doubts on whether this motion is specific to quasicrystals or just a generic feature of systems that contain structural vacancies, and whether phason jumps are qualitatively different, aside from their low activation energy, in comparison to single vacancy diffusion. Additional evidence against phason-mediated diffusion has also been reported [59].
From the foregoing discussion, we may expect the dynamics of quasicrystals to be distinct in some ways from the dynamics of crystalline solids, and some similarities to glass-forming liquids can be anticipated. Molecular dynamics (MD) simulations provide a suitable tool for elucidating basic thermodynamic and dynamic properties of quasicrystals. In this regard, Engel and Trebin [60] introduced the Lennard-Jones-Gauss (LJG) system which constitutes a simple model for studying quasicrystals. The potential has proven to be able to reproduce several entropically stabilized quasicrystal phases, e.g., see Ref. 61. From an atomistic level, MD simulations also provided early evidence for short-time stochastic particle dynamics in the form of phason-flips [62]. Interestingly, this can happen either with single-particle jumps but also, and more interestingly, with correlated ring-like multi-particle motion [62] or carousel-like dynamics [63]. This string-like collective dynamics is evidently another commonality between quasicrystals and glasses.
Despite the strong ongoing interest in quasicrystals from different directions, a complete description and understanding of their vibrational, thermodynamic and collective properties is absent. Moreover, a final answer to the question of whether phasons exist and which is their microscopic origin have not been reached yet. Finally, the understanding of quasicrystal dynamics from an atomistic perspective and the precise characterization of their degree of similarity with amorphous solids and liquids are still largely unexplored. The present work aims to provide a first step towards this direction by carrying out an exhaustive analysis of a 2D dodecagonal quasicrystal using MD simulations.
Figure 2: The LJG-type potential and phase diagram. (a) Double-well nature of the potential with a Gaussian dip. (b) Two phase transitions upon cooling. (c)-(e) Atomic configurations of quasicrystal, hexatic phase and liquid at temperatures $T=0.3$, $1.2$ and $1.95$, respectively.
In the present study, we delve into a comprehensive examination of the thermodynamic properties, dynamic characteristics, and the structural origins of dynamics in quasicrystals. Our analysis is based on atomistic simulations, focusing on a model system comprising identical particles interacting in two dimensions through the LJG-type pair potential. Sec. II is dedicated to introducing the LJG-type potential, outlining the phase diagram, describing the dodecagonal quasicrystal system, and providing essential details regarding our computer simulations. In this section, we consider an exploration of the thermodynamics of a dodecagonal quasicrystal. In Sec. III, we delve into the study of relaxation mechanisms, dynamic heterogeneity, bulk self-diffusion, and the intricate interplay among these properties. At lower temperatures, we observe phason flips as part of the dynamic behavior in quasicrystals. Meanwhile, at higher temperatures, our analysis reveals temperature-dependent string-like motions within the quasicrystal. Additionally, we uncover that both individual phason flips and collective string-like motions originate from the transfer of defects within the quasicrystal. Furthermore, we introduce the concept of “string lifetime” as a important characteristic timescale in the dynamics of quasicrystals. Finally, in Sec. IV, we conclude with a comprehensive discussion of our findings. Our research offers valuable insights into the fundamental physical and mechanistic aspects governing the thermodynamic and dynamic properties of dodecagonal quasicrystals and other quasicrystalline structures.
II Methods
II.1 Lennard-Jones-Gauss potential
We conducted classical molecular dynamics (MD) simulations using the LAMMPS (Large-scale Atomic/Molecular Massively Parallel Simulation) software [64] to investigate the thermodynamic and dynamic properties of a 2D dodecagonal quasicrystal. Our system comprises identical particles interacting via a novel LJG-type potential, which combines the traditional Lennard-Jones potential with a secondary Gaussian well at longer distances. The LJG potential is defined as follows:
$V\left(r\right)=4{\varepsilon_{0}}\left[{{{\left({\frac{r}{{{\sigma_{0}}}}}\right)}^{-12}}-{{\left({\frac{r}{{{\sigma_{0}}}}}\right)}^{-6}}}\right]+\frac{h}{{{\sigma{\rm{G}}}\sqrt{2\pi}}}\exp\left[{-\frac{{{{\left({r-{r_{\rm{G}}}}\right)}^{2}}}}{{2\sigma_{\rm{G}}^{2}}}}\right].$ (1)
In this equation, $\varepsilon_{0}$ represents the energy unit, and $\sigma_{0}$ is the length unit. The parameter $h$, in conjunction with the standard deviation $\sigma_{\text{G}}$, dictates the peak height of the Gaussian component, while $r_{\mathrm{G}}$ denotes the peak position of the Gaussian well. After extensive experimentation, the chosen parameter values that resulted in the formation of a dodecagonal quasicrystal are $\varepsilon_{\mathrm{0}}=0.5$, $\sigma_{\mathrm{0}}=0.88$, $h=-0.874$, $r_{\mathrm{G}}=1.879$, and $\sigma_{\textrm{G}}^{2}=0.02$. All quantities are expressed in dimensionless Lennard-Jones units. Notably, with these carefully selected parameters, the LJG-type potential exhibits a double-well nature, as illustrated in Fig. 2(a). For practical purposes, we applied a cutoff to the LJG-type potential at $r=2.5$.
II.2 Phase diagram
A system comprising $N=1024$ particles arranged in a square lattice is subjected to a controlled heating process, transitioning into the liquid phase at a temperature of $T=1.95$. The simulations are conducted with periodic boundary conditions applied in both spatial dimensions. Following an initial period of 50,000 molecular dynamics (MD) steps to ensure equilibration, the system undergoes a controlled linear temperature decrease to reach $T=0.001$ over a course of $3.9\times 10^{7}$ MD steps. To maintain precise control over temperature and pressure within the isothermal-isobaric ensemble, we employ two crucial techniques. The Nosé-Hoover thermostat, as introduced by Nosé [65] and further refined by Hoover [66], regulates temperature fluctuations throughout the simulations. Meanwhile, the Parrinello-Rahman barostat, developed by Parrinello and Rahman [67], ensures stable pressure conditions.
During the cooling process, the system exhibits two first-order phase transitions. The transition from phase 1 (liquid) to phase 2 (hexatic) happens at about $T_{LH}=1.78$. In order to quantitatively identify the location of the transition between these phases, and qualify the nature of phase 2, we calculate the sixfold bond-orientation order parameter [68, 69] according to
$\psi_{6}^{j}=\frac{1}{N_{j}}\sum_{m=1}^{N_{j}}\exp\left(i6\theta_{m}^{j}\right),$ (2)
where $N_{j}$ is the number of nearest neighbors of particle $j$, $\mathbf{r}_{j}$ is the position of particle $j$, and $\theta_{m}^{j}$ is the angle between the vector ($\mathbf{r}_{m}-\mathbf{r}_{j}$) and the $x$ axis. The cutoff distance for determining these nearest neighbors is set to be $1.2$, which corresponds to the bond length derived from the radial distribution function (RDF), $g(r)$, as discussed below. The average order parameter is then obtained via $\overline{\Psi}_{6}=\sum_{j=1}^{N}\psi_{6}^{j}$. The temperature dependence of $\hat{\psi}_{6}$ is shown in Fig. 3(a), which clearly indicates two thermodynamic transitions at high and intermediate temperatures.
Furthermore, the spatial correlation of the particle-level $\psi_{6}^{j}$ can be computed using the following formula:
$g_{6}(r)=\frac{L^{2}}{2\pi r\Delta rN\left(N-1\right)}\left\langle\sum_{j\neq k}\delta\left(r-\left|\mathbf{r}_{jk}\right|\right)\psi_{6}^{j}\psi_{6}^{k\ast}\right\rangle,$ (3)
where $L$ is the length of the simulation box, and $N$ the number of atoms. The quasi-long-range bond-orientation order arises from the algebraic decay of $g_{6}\left(r\right)/g\left(r\right)\thicksim r^{-\eta_{6}}$. It is noteworthy that at the boundary between the liquid phase and the hexatic phase, $\eta_{6}$ assumes a value of 1/4. In the liquid phase, the system exhibits only short-range bond-orientation order. Consequently, the bond-orientation correlation function is more aptly described by the Ornstein-Zernike (OZ) function, characterized by $g_{6}\left(r\right)/g\left(r\right)\thicksim r^{-1/2}\exp\left(-r/\xi_{6}\right)$ [70, 71, 68]. As evident in Fig. 3(b), $g_{6}(r)/g(r)$ exhibits a gradual decay with increasing temperature. We observe a power-law decay with an exponent of $\eta_{6}=-1/4$ for $g_{6}(r)/g(r)$ approximately at $T_{LH}=1.78$, signifying the location of the liquid-hexatic transition.
Figure 3: (a) The temperature-dependent sixfold bond orientation order parameter. (b) Comparison of $g_{6}(r)/g(r)$ at various temperatures in the vicinity of $T_{LH}=1.78$. The solid lines represent the outcomes of OZ fittings, while the dashed lines depict the results of power-law fittings. The green dashed line marks the boundary at $g_{6}(r)/g(r)\thicksim r^{-1/4}$, signifying the demarcation between the liquid phase and the hexatic phase.
The transition from phase 2 (hexatic) to phase 3 (quasicrystal) takes place at approximately $T_{HQ}=0.87$. Through a detailed comparison of the local configurations in phase 3 with experimental observations obtained via high-resolution transmission electron microscopy (TEM) [72], as illustrated in Fig. 4, we confirm that phase 3 corresponds to a monolayer dodecagonal (12-fold) symmetry quasicrystal, pointing to the relevance of our simulations to real-world quasicrystal material. Our model quasicrystal exhibits dodecagonal units and a point defect structure within a hexagonal ring configuration, both of which align with the observations made through TEM [72] and other experimental observation [26]. The agreement between simulation results and experimental observations confirms the fidelity of our model in capturing the intricate atomic arrangement of this remarkable quasicrystal phase. Further exploration of this configuration, both in direct and reciprocal space, will be the focus of our analysis in Section III.
We note that there have been many previous simulations of two dimensional crystalline materials in which corresponding transitions from a low temperature hexagonal crystalline state to an orientationally correlated hexatic state occurs upon heating, followed by an additional transition between the hexatic and liquid states at an even higher characteristic temperature. Although the well-known theories of these transitions by Kosterloitz and Thouless [73] and Halperin and Nelson [74, 75], many studies have established that these transitions can be either of second or first order in nature [76], depending on the particular form of the interparticle potential, and the presence of appreciable finite size effects can often make the assignment of transition order computationally challenging and sometimes controversial. Resolving the difference between weakly first order transitions and second order transitions can be inherently difficult [77].
There have also been studies of the collective dynamics of two dimensional systemes exhibiting crystal melting and formation of an intermediate hexatic phase [78, 79, 80]. In the present exploratory study of a model quasi-crystalline material we simply note that the thermodynamic transitions that we observe in our quasi-crystal material simulations are strikingly reminiscent to the hexagonal to hexatic and hexatic to liquid transitions established to exist in many two-dimensional crystal forming materials. Moreover, these transitions in our quasi-crystalline forming material are relatively sharp, suggestive of a weakly first order phase transitions. Given the limited size of our simulated materials, however, we do not insist on this designation of transition order and simply note that apparently well-defined thermodynamic transitions occur near certain characteristic temperatures at which the degree of material order and dynamics clearly occur. We are not aware of a previous work exploring these transitions in quasi-crystalline materials.
Previous work on crystal melting have also found a prevalence for string-like collective motion [78, 79], a phenomenon that appears superficially very similar to the observations that we report below in our simulated model quasi-crystal forming material. Despite the superficial similarities in the thermodynamic and dynamic properties of two dimensional crystal and quasi-crystal forming materials, we stop short of equating their properties. Nonetheless, the evidence of evident long-range positional correlations in our quasicrystalline material point to a greater ”similarity” to crystalline materials than glass-forming materials. However, given the somewhat contradictory indications from simulation and measurements of a ”similarity” of quasi-crystalline materials to both crystalline and glass-forming materials, we take a fresh look at which class of materials more closely resembles quasi-crystalline materials.
Figure 4: The atomic structure of the monolayer dodecagonal (12-fold) quasicrystal, following relaxation at $T=0.3$, finds support from TEM observations [72].
II.3 Quasicrystal structure and dynamics
On one hand, the radial distribution function (RDF) $g(r)$ provides insights into the probability of locating a particle at a distance $r$ from the reference particle’s position in direct space. In the case of crystalline materials, $g(r)$ exhibits distinct peaks corresponding to the periodic arrangements of particles, signifying the presence of long-range order. For isotropic systems, $g(r)$ relies solely on the magnitude of the relative atomic distances and can be calculated as follows:
$g\left(r\right)=\frac{1}{2\pi\,r\,dr}\frac{1}{\rho}\left\langle\sum_{i\neq i^{\prime}}\delta\left(r-\lvert\mathbf{r}_{i}-\mathbf{r}_{i^{\prime}}\rvert\right)\right\rangle,$ (4)
where $\rho$ represents the number density.
On the other hand, the static structure factor $S(k)$ is another essential parameter for describing the local structure of materials. It characterizes the system’s structure in reciprocal space and can be directly measured in experiments involving elastic scattering (diffraction). In MD simulations, $S(k)$ is computed directly from atomic positions based on the relation,
$S\left(k\right)=\frac{1}{N}\left\langle\sum_{j,j^{\prime}=1}^{N}\exp\left\{-i\mathbf{k}\cdot\left[\mathbf{r}_{j}-\mathbf{r}_{j^{\prime}}\right]\right\}\right\rangle.$ (5)
These two quantities, $g\left(r\right)$ and $S\left(k\right)$, play pivotal roles in characterizing the structure of phases from two complementary perspectives.
Furthermore, the mean-squared displacement (MSD) serves as a standard method to probe the dynamics of materials. It typically quantifies the average displacement of a particle within the system as a function of time, expressed as,
$\left\langle r^{2}\left(t\right)\right\rangle=\frac{1}{N}\left\langle\sum_{i=1}^{N}{\lvert\mathbf{r}_{i}\left(t\right)-\mathbf{r}_{i}\left(0\right)\rvert}^{2}\right\rangle.$ (6)
However, in 1D and 2D systems, long-range thermal fluctuations in positional ordering, as described by the Mermin-Wagner fluctuations [81], can pose challenges for accurately measuring real dynamics. These fluctuations manifest as appreciable effects, as the system size imposes a cutoff on the maximum wavelength [82, 83, 84].
[…]
Mitigating the finite-size effects when studying dynamics in the 2D dodecagonal quasicrystal system is especially important when using periodic boundary conditions. To address this problem, we also employ a modified method by calculating the “cage-relative MSD” (CR-MSD) in place of the standard MSD:
$\left\langle r^{2}\left(t\right)\right\rangle_{\rm{CR}}=\frac{1}{N}\left\langle\sum_{i=1}^{N}\left[\left(\mathbf{r}_{i}\left(t\right)-\mathbf{r}_{i}\left(0\right)\right)-\frac{1}{N_{j}}\sum_{j=1}^{N_{j}}\left(\mathbf{r}_{j}\left(t\right)-\mathbf{r}_{j}\left(0\right)\right)\right]^{2}\right\rangle.$ (7)
In this context, the second term in Eq. (7) represents the center of mass of the cage, which is derived from the positions of the $N_{j}$ nearest neighbors of the central particle $i$. This approach helps mitigating finite-size effects and allows for a more accurate assessment of dynamics in the 2D dodecagonal quasicrystal system.
In most cases, particles exhibit ballistic motion at very short times and eventually transition to random diffusive motion at longer timescales. As such, Mean square displacement (MSD), or the cage-relative MSD (CR-MSD in the context of a 2D system), can be employed to assess the diffusivity $D$ of a $d$-dimensional system, utilizing the Einstein relation:
$D=\frac{1}{2d}\lim_{t\to\infty}\frac{\left\langle r^{2}\left(t\right)\right\rangle_{\rm{CR}}}{t}.$ (8)
The determination of this quantity provides an important point of contact between experiments on quasi-crystalline materials and our simulations.
In addition to MSD, another observable property that quantifies dynamic relaxation is the self-intermediate scattering function (SISF), which serves to characterize density fluctuations of the same particle within the system at time $t=0$ and at a subsequent time $t$. This quantity is defined as,
$F_{s}\left(k,t\right)=\frac{1}{N}\left\langle\sum_{i=1}^{N}\exp\left\{-i\mathbf{k}\cdot\Delta\mathbf{r}_{i}\left(t\right)\right\}\right\rangle,$ (9)
where $\Delta\mathbf{r}_{i}\left(t\right)=\mathbf{r}_{i}\left(t+\tau\right)-\mathbf{r}_{i}\left(\tau\right)$, with $\tau$ representing a time interval. The wavenumber $k$ is chosen as the first peak of the static structure factor $S(k)$, which in the present system this corresponds to $k^{*}$. The structural relaxation time $\tau_{\alpha}$ is then defined as the time when $F_{\text{s}}\left(k^{*},\tau{\alpha}\right)=e^{-1}$.
Similar to the approach used with CR-MSD, we can define a cage-relative intermediate scattering function based on Eq. (9). Here, we subtract the motion of the center of mass of the nearest neighbors from the total displacement:
$F_{s\_\rm{CR}}\left(k^{*},t\right)=\frac{1}{N}\left\langle\sum_{i=1}^{N}\exp\left\{-i\mathbf{k}\cdot\left(\Delta\mathbf{r}_{i}\left(t\right)-\Delta\mathbf{r}_{i\_\rm{CR}}\left(t\right)\right)\right\}\right\rangle,$ (10)
where the displacement of the center of mass given by $N_{i}$ neighbors reads
$\Delta\mathbf{r}_{i\_\rm{CR}}\left(t\right)=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\left(\mathbf{r}_{j}\left(t+\tau\right)-\mathbf{r}_{j}\left(\tau\right)\right).$ (11)
This approach allows us to account for nearest neighbor contributions and effectively characterize the dynamics in the presence of structural constraints.
The non-Gaussian parameter $\alpha_{2}\left(t\right)$ stands as a fundamental measure of dynamic heterogeneity in complex systems, indicating the degree of deviation from a Gaussian distribution in the displacement of all particles. In 2D systems, $\alpha_{2}\left(t\right)$ can be calculated as follows:
$\alpha_{2}\left(t\right)=\frac{\left\langle\sum_{i=1}^{N}\left[\mathbf{r}_{i}\left(t\right)-\mathbf{r}_{i}\left(0\right)\right]^{4}\right\rangle}{2\left\langle\sum_{i=1}^{N}\left[\mathbf{r}_{i}\left(t\right)-\mathbf{r}_{i}\left(0\right)\right]^{2}\right\rangle^{2}}-1.$ (12)
However, in line with the approach employed for the cage-relative MSD and SISF, we precisely calculate the cage-relative non-Gaussian parameter, which is defined as:
$\alpha_{2\_\rm{CR}}\left(t\right)=\frac{\left\langle\sum_{i=1}^{N}\left[\left(\mathbf{r}_{i}\left(t\right)-\mathbf{r}_{i}\left(0\right)\right)-\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\left(\mathbf{r}_{j}\left(t\right)-\mathbf{r}_{j}\left(0\right)\right)\right]^{4}\right\rangle}{2\left\langle\sum_{i=1}^{N}\left[\left(\mathbf{r}_{i}\left(t\right)-\mathbf{r}_{i}\left(0\right)\right)-\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\left(\mathbf{r}_{j}\left(t\right)-\mathbf{r}_{j}\left(0\right)\right)\right]^{2}\right\rangle^{2}}-1.$ (13)
This refined approach accounts for the motion of the center of mass of the nearest neighbors and provides a more precise assessment of the non-Gaussian behavior in 2D system.
Figure 5: Temperature dependence of the density, highlighting two distinct phase transitions as the system undergoes cooling, which link the liquid phase, hexatic phase, and quasicrystal phase.
III Results
III.1 Thermodynamic properties and structure
III.1.1 Density
During the rapid cooling process from the high-temperature liquid phase, the density displays two distinct transitions, as illustrated in Fig. 5. The first transition occurs at approximately $T_{LH}=1.78$, marking the shift from the liquid phase to the orientationally correlated hexatic phase. Subsequently, the second transition takes place as the system undergoes a transformation from the hexatic phase to a dodecagonal quasicrystal phase, occurring at around $T=0.87$. It is worth noting that the transition from the hexatic to the quasicrystal phase is characterized by a positive thermal expansion. In contrast, during the transition from the liquid phase to the hexatic phase, there is no observable expansion or compression, as depicted in Fig. 5.
III.1.2 Heat capacity
We calculate the heat capacity using the ”direct” method of taking the derivative of the total energy with respect to temperature in cooling paths. Although this method, while applicable in non-equilibrium heating/cooling paths, may large uncertainties in estimating thermodynamic properties in equilibrium [86]. It is sufficient for calculating the relative value of different phases and representing phase transitions. In a manner similar to the density, the heat capacity as a function of temperature (Fig. 6) reveals two well-defined thermodynamic transitions. These transition temperatures ($T_{LH}=1.78$ and $T_{HQ}=0.87$) are consistent with previous findings. The two transitions seem sharp, but it is not possible to deduce the transition order due to the limited system size as above mentioned.
Figure 6: Temperature dependence of heat capacity, revealing two prominent phase transitions that correlate with shifts in density, as depicted in Fig. 5.
III.1.3 Radial distribution function
The RDF is a widely utilized tool for elucidating the structural characteristics of ordered and disordered materials. In Fig. 7, we present the RDF for the liquid phase at $T=1.95$, the hexatic phase at $T=1.2$, and the dodecagonal quasicrystal at $T=0.3$. The RDF of the quasicrystal exhibits a distinctive behavior that positions it between the characteristics of crystalline and disordered materials. This observation supports the notion of an intermediate state of matter, as alluded to in the title and introduction of our paper.
In addition to the well-established RDF profiles in amorphous and hexatic phases, the quasicrystal exhibits unique features. In the short range, the RDF of the quasicrystal reveals distinct peaks, which can be attributed to the quasi-periodic tiling. The first peak emerges at approximately $r=1.0$, while the second peak appears around $r=1.89$. These values are readily explicable as they correspond to the positions of potential energy wells (stable states) within the LJG-type potential (refer to Fig. 2(a)). Conversely, at longer distances, the RDF of the quasicrystal displays typical characteristics of disordered materials. It gradually converges towards a constant value with non-vanishing fluctuations, reflecting the absence of long-range translational symmetry, akin to amorphous systems, while still retaining certain translational symmetry.
Figure 7: Radial distribution function of (a) liquid phase at $T=1.95$, (b) hexatic phase at $T=1.2$ and (c) dodecagonal quasicrystal at $T=0.3$.
III.1.4 Static structure factor
In MD simulations, a common method for computing $S(k)$ is through the Fourier transform of the RDF $g(r)$ [87]. Specifically, $S(k)$ is given by the equation:
[…]
where $\rho$ is the number density of particles, and $k$ is the wave vector.
As depicted in Fig. 8, the static structure factors of the liquid, hexatic, and quasicrystal phases all prominently feature a main peak around $k^{*}=2\pi/a$, where $a$ represents the nearest-neighbor distance in real space. A pronounced peak in the structure factor at $k^{*}$ serves as a indication of local correlations in the particle positions in the condensed material. This peak corresponds to a wave vector that generates a long-range oscillatory behavior with a periodicity of $a$ in real space [88]. Moreover, the height of the first peak in the static structure factor is a critical parameter for predicting the liquid-solid phase transition. According to the Hansen-Verlet criterion in crystalline materials the freezing transition occurs when the value of the first peak in $S(k)$ reaches a specific constant value [89]. In the liquid phase, the first peak is notably lower than those in the hexatic and quasicrystal phases, underscoring the lack of long-range order and the inherent disorder in the system [90].
Figure 8: Static structure factor of the (a) liquid phase at $T=1.95$, (b) hexatic phase at $T=1.2$ and (c) dodecagonal quasicrystal at $T=0.3$.
Figure 9: Intensity map of the static structure factor of (a) liquid phase at $T=1.95$, (b) hexatic phase at $T=1.2$ and (c) quasicrystal phase at $T=0.3$.
We have also generated the intensity maps of the static structure factor using Eq. (5), which provide a comprehensive representation of the 2D structures included in the phase diagram. These intensity maps resemble the non-energy-resolved diffraction patterns typically obtained in diffraction experiments to determine crystal structure and symmetry. The diffraction pattern of the disordered phase (Fig. 9(a)) consists of diffuse rings of light, resulting from the lack of regular arrangement in ordered lattices. In contrast, the diffraction pattern of the hexatic phase exhibits both weak polyhedra and bright reflections with 6-fold and 12-fold symmetries (Fig. 9(b)), clearly indicating the hexatic phase intermediate nature between the liquid and solid phases. The diffraction pattern of the quasicrystal (Fig. 9(b)) is more intricate, displaying a 12-fold symmetry and well-defined reflections arranged in a quasi-periodic 2D pattern. These diffraction patterns serve to reinforce the credibility of our atomistic model. Note that for each of these diffraction patterns, we conducted parallel simulations on 50 independent configurations to enhance statistical reliability.
III.2 Dynamic properties
III.2.1 Structural relaxation
In this section, we explore the SISF and the structural relaxation or $\alpha$-relaxation time of the dodecagonal quasicrystal. Fig. 10(a) illustrates the temporal evolution of CR-$F_{s}(k^{*},t)$ as the temperature increases, calculated using Eq. (10). Here, the SISF is defined at the wave vector $k=7.3$, which corresponds to the major peak in $S(k)$ and is sensitive to variations in atomic density, making it an important indicator of structural transitions, particularly in disordered phases. SISF focuses on the timescale for the decay of density-density correlations of matter, and it is typically defined at the first peak of $S(k)$. The evolution of the SISF reveals a two-step relaxation process, including the short-time $\beta$ relaxation and the long-time $\alpha$-relaxation. The $\alpha$-relaxation time $\tau_{\alpha}$ is defined as the time when $F_{s}(k^{*},\tau_{\alpha})=1/e$, as indicated by the horizontal dashed line in Fig. 10(a).
Figure 10: Relaxation dynamics in the quasicrystal. (a) Cage-relative SISF vs time at various temperatures ranging from $T=0.364$ to $T=0.825$. (b) Fitted curves of the relaxation function using Eq. (15) to the cage-relative SISF data. (c) Arrhenius plot illustrating the temperature-dependent $\alpha$-relaxation time, with a dashed line indicating the linear fit obtained from high-temperature data. The inset shows the temperature dependence of the activation free energy $\Delta G(T)$.
It is evident that structural relaxation, as measured by the self-intermediate scattering function, occurs in our quasicrystalline material, is similar to both glass-forming liquids [91] and heated crystalline solids [85] in the sense that there is a general fast beta relaxation occurring on a timescale on the order of a ps that is relatively temperature insensitive and a longer time alpha relaxation process having a relaxation time that is highly temperature dependent.
As the temperature decreases, the $\alpha$-relaxation time increases rapidly, a well-known, even defining, attribute to glass-forming liquid, but this behavior is also characteristic to heated crystals and derives from the reduced kinetic of the atoms in the crystal and the associated reduced rate of atomic diffusion [85, 85]. Eventually, at a certain crossover temperature in the liquid regime (not study in the present paper), the $\alpha$- and fast $\beta$-relaxation processes merge, defining an onset temperature $T_{A}$ for the multi-step dynamics exhibited by glass-forming, crystalline and quasi-crystal materials. This pattern of relaxation thus seem to be a universal pattern of relaxation in condensed matter rather broadly. At low temperatures, the dynamics of the quasicrystal becomes so slow that it becomes increasingly difficult to make simulation observations under equilibrium conditions, a familiar situation in glass-forming liquids, but again this is the case in crystalline materials [92]. To quantity the observations in Fig. 10, a necessary step in comparing to the dynamics of crystalline and glass-forming materials sufficiently to enable determination which material is more like our quasi-crystalline system, we fit our $F_{s}(k,t)$ data to a two-step relaxation function, defined as,
$F_{s}\left(k^{*},t\right)=\left(1-A_{\alpha}\right)e^{-\left(t/\tau_{f}\right)^{\beta_{f}}}+A_{\alpha}e^{-\left(t/\tau_{\alpha}\right)^{\beta}}.$ (15)
This function has been found to describe $F_{s}\left(k^{*},t\right)$ for a range of materials [93, 94].
Our fitting of the fast dynamics properties $\tau_{f}$ and $\beta_{f}$ leads to $\tau_{f}\approx 0.1$ and $\beta_{f}\approx 1.9$ (Fig. 10(b)), which is rather typical of glass-forming liquids [95]. The ”mode amplitude” or ”non-ergodicity” parameter $A_{\alpha}$ displays relatively weak temperature dependence within the explored temperature range. Importantly, fixing these three parameters ($\tau_{f}$, $\beta_{f}$, and $A_{\alpha}$) does not significantly compromise the accuracy of our fits across the entire temperature range investigated[94], allowing for precise quantification of our quasi-crystal relaxation. At this stage, it is not evidently clear whether our quasi-crystal material is more like a ”typical” crystalline material or a glass-forming liquid as these materials share many features in their dynamics when temperature is varied. We next consider the temperature dependence of $\tau_{\alpha}$ to search for some discriminating characteristics of the quasi-crystal dynamics that could help us in this determination.
In Fig. 10(c), we present the $T$ dependence of $\tau_{\alpha}$ in our 2D dodecagonal quasicrystal in the traditional form of an Arrhenius plot. We observe a relatively high temperature regime in which $\log\tau_{\alpha}$ exhibits roughly linear variation with $1/T$, but the slope bends over sharply as the $T$ decreases below 0.5, a temperature notably in the quasicrystal phase regime. We emphasis here that we did not simulate our quasi-crystalline material in the liquid regime so that should be another Arrhenius-like regime for $\tau_{\alpha}$ at higher temperature, with probably a different slope (activation energy). We thus simulate an intermediate temperature range where multi-step relaxation regime and cannot assign the onset temperature of glass-forming corresponding to a temperature $T_{A}$ where the $\alpha$ and fast $\beta$ relaxation processes start to merge, and where relaxation can be reliably be taken as being Arrhenius and non-cooperative to a good approximation [96, 97]. Qualitatively, the kink-like transition from Arrhenius to non-Arrhenius behavior aligns with numerous previous observations on some glass-forming materials [98, 99], but also highly anharmonic crystalline materials, such as superionic crystalline materials [100]. (We will discuss this Arrhenius curve ”kink” further below in relation to its genreal occurrence in the intermediate forms of matter mentioned by Angell [10].)Thus, we still have not arrived at the desired descrimination of quasi-crystalline materials in relation to other more studied condensed materials.
Past experience on both glass-forming liquids [95] and heated crystalline materials [100] has shown the estimation of the activation energy from the local slope of this type of Arrhenius plot cannot be reliably used to define the activation energy under conditions in which the slope of the Arrhenius curve varies with $T$ can be highly misleading [95]. An estimate of the activation free energy for diffusion requires that the actual $T$ dependent activation free energy $\Delta G(T)$, formally defined by the relation, $\Delta G(T)\equiv k_{B}T\ln(\tau_{\alpha}/\tau_{0})$ [97, 101, 95]. Here, $\tau_{0}$ serves as a prefactor associated with the vibrational relaxation time, estimated at approximately 0.05 in our reduced units. Estimates of the temperature dependent activation energy deduced from this definition are indicated in the inset of Fig. 10(c) where we observe that $\Delta G(T)$ progressively increases upon cooling over the temperature range that we investigate. We interpret this trend to be consistent with the sigmoidal temperature variation of $\Delta G(T)$ observed in previous studies of structural relaxation of Al-Sm, a glass-forming liquid [97, 95] that exhibits fragile to strong type glass-formation, and experiments and simulation on superionic crystalline materials [100]. Again, we have not been able to resolve whether the dynamics is more like a crystal or a heated crystalline system, forcing us the nature of dynamic heterogeneity in the quasi-crystal material in the hopes of finally resolving the situation.
III.2.2 Dynamic heterogeneity
To capture the “dynamic heterogeneity”, i.e., mobility fluctuations of the dodecagonal quasicrystal, we employ the non-Gaussian parameter frequently used in the literature, $\alpha_{2}\left(t\right)$ [102, 103]. This parameter quantifies the degree of deviation of particle displacements from a Gaussian distribution and this phenomenon has often been found in association with large mobility fluctuations in complex liquids, making this quantity an indirect measure of such dynamic heterogeneity. The time evolution of CR-$\alpha_{2}\left(t\right)$ for the dodecagonal quasicrystal is depicted in Fig. 11(a) at various temperatures. Similar to glass-forming liquids, $\alpha_{2}(t)$ exhibits a peak at a characteristic time $t^{*}$ [102] that is generally shorter than $\tau_{\alpha}$. This measure of the ”lifetime” of the dynamic heterogeneity increases upon decreasing temperature, a trend typical of glass-forming liquids, but contrary to heated crystals where this quantity appears to increase upon heating from the low temperature crystal state where this quantity is very small, but then passes through a maximum at intermediate temperatures and then falls upon approaching the melting temperature[85]. This decreases of $\alpha_{2}$ upon cooling was found to be especially prevalent in the simulation of superionic ${\rm UO}_{2}$ in its high temperature [100]. This is the first hint that the dynamics of quasi-crystalline material is perhaps more like a glass-forming liquid at low $T$ than a heated crystal. However, we do not consider the $T$ regime at which the hexatic phase of the quasicrystalline material enters the liquid states, the analog of the melting in the 3-dimensional crystalline material studied previously. In contrast, we find that $\alpha_{2,\text{max}}$ progressively increases upon lowering $T$, a pattern of behavior that aligns with the dynamics of glass-forming liquids rather than crystalline materials [104]. This trend clearly points to a commonality between glass-forming liquids and our quasi-crystalline material. We next consider the interrelation between diffusion and the structural relaxation time and $t^{*}$, which as we shall see offers further and more compelling evidence for a closer relationship between the dynamics of our quasi-crystalline material and glass-forming material rather than the dynamics of crystalline materials.
Figure 11: (a) Cage-relative non-Gaussian parameter $\alpha_{2}$ as a function of time at different temperatures. (b) Temperature-dependence of the maximum value of the CR-$\alpha_{2}$, i.e., $\alpha_{2,\text{max}}$. We expect the linear variation in this figure to only hold over a limited $T$ range.
III.2.3 Revisiting the kink in the Arrhenius curve of relaxation
In Fig. 12(a), we show the evolution of the temperature dependence of CR-MSD for the dodecagonal quasicrystal when the temperature is varied over a large range. As in condensed material rather generally, particle MSD exhibits three stages that mirrors these same dynamical regimes in $F_{s}(k^{*},t)$. At short times, particles undergo ballistic motions, i.e., a particle does not collide with any particles at very short times so that the motion resembles particles in a gas. During this stage, the distance traveled varies nearly linearly with time, resulting in MSD being proportional to $t^{2}$. Beyond this ”inertial dynamics” regime, a plateau in $\left\langle\Delta r^{2}\left(t\right)\right\rangle$ develops which becomes more pronounced at lower temperatures. This pattern of behavior is characteristic of condensed materials rather broadly so it is no surprise we see the same basic trend in our quasi-crystal material. At long times, the particles escape from their ”cages” at intermediate time in which most of the atoms vibrate within regions defined by surrounding particles. The Debye-Waller parameter $\left\langle u^{2}\right\rangle$, which describes the amplitude of atomic vibrations on the fast $\beta$ relaxation time (see Fig. 10(a)), defining the onset of caging and the ”end” of the fast $\beta$ relaxation process. At sufficiently long times, particle motion can be described as simple diffusion, provided the material is in thermal equilibrium. Of course, the timescale required for this long time diffusion emerges, corresponding to a time on the order of $\tau_{\alpha}$ can be extremely large and difficult to access by molecular dynamics simulation. This can make the estimation of the diffusion derived from particle displacement difficult to estimate computationally as we shall next show.
Figure 12: (a) Cage-relative MSD at different temperatures. (b) Fit of MSD to the von Schweidler expression at two representative temperatures. (c) Arrhenius plot of diffusivity, where a striking transition of mechanism is found at $T=0.44$. At high temperatures (hollow points), $D_{0}=1.07\times 10^{-4}$ and $\Delta E=7.38$, while for low-temperature regime (solid points), $D_{0}=6.07\times 10^{-10}$ and $\Delta E=1.54$.
As in glass-forming, we utilize the semi-empirical expression [102, 105] to estimate $D$ from the widely utilized relation on this context,
$\left\langle\Delta r^{2}\left(t\right)\right\rangle=r_{0}^{2}+a\left(4Dt\right)^{b}+4Dt.$ (16)
While this method of estimating $D$ provides a reasonable method for estimating $D$ in our quasi-crystalline material, the slowing of the dynamics limits the temperature range in which estimates can be made and we confine our estimates to the same $T$ range as we considered above for $\tau_{\alpha}$. Fig. 12(b) shows the result of our fitting procedure based on Eq. (16) to estimate $D$ as function of $T$.
As in the case of $\tau_{\alpha}$, we see that $D$ exhibits two distinct temperature regimes when the data is considered in the form of an Arrhenius plot, as depicted in Fig. 12(c). Measurements of $D$ in a real quasi-crystalline material have indicated a striking similar ”kink” in the Arrhenius curve for $D$ [50], confirming this interesting feature found in our simulations. Recent simulations of $D$ for a model metallic glass material [95] also strikingly resemble our observations in Fig. 12, where a kink change arises in the Arrhenius curve for $D$ data indicating a clear change in the dynamical regimes of the material where the origin of this change is investigated in depth [95, 101]. Arrhenius curves having a ”kink” as in Fig. 12 having upward curvature characteristically arise in the dynamics of enzymes [106, 107, 108], an observation in line with Angell’s suggestion that proteins could also be viewed as an intermediate form of matter having properties in common with other materials of this ypte. It was mentioned before that this type of ”kink” in the Arrhenius curves is found in Al-Sm metallic glass-forming liquids [95, 101] and moreover this type of non-Arrhenius dynamics is observed in simulations and measurements of the dynamics of model superionic materials, and the interfacial dynamics of crystalline materials [109], providing further evidence supporting Angell heuristic arguments.
Below we show that the temperature variation of $D$ and $\tau_{\alpha}$ bares a direct relation to $t^{*}$ and $\tau_{\alpha}$ so we do not provide a discussion of the $T$ dependent activation energy as it must be the same as deduced from Fig. 12(c). However, we do point out, contrary to the visual impression of the Arrhenius plots of Fig. 10(c) and Fig. 12(c), that the activation energy monotonically increases upon cooling rather than decreasing upon cooling.
Figure 13: Characterization of the atomic motion mechanisms in quasicrystal by displacement vector (yellow arrows) at different temperatures. The emergence of collective 1D string-like structures is evident at high temperatures.
III.2.4 Atomic-scale mechanisms governing relaxation and diffusion in the fast $\beta$ and $\alpha$ relaxation regimes
To further investigate the mechanisms of the transition between a distinct fast $\beta$ dynamics regime and $\alpha$relaxation and the curious kinks in the Arrhenius curves for $\tau_{\alpha}$ and $D$, based on a direct examination on the shape of particle trajectories and the occurrence of collective motion in which more than one particle is involved. First, we examine the atomic trajectories and correlated motion at different temperatures over $3\times 10^{7}$ MD steps. At the lowest $T$ considered, the most prevalently observed atomic displacements, apart from relatively small-scale vibrations, involve isolated local particle hopping, that might be compatible with hypothetical “phason flips” suggested to exist in quasicrystal systems [110]. These ”defects” transform energetically comparable configurations into others and remind us of moving interstitials in crystalline materials. Larger scale collective motions might possibly at these low temperatures, but if they do occur, they must be rare. Starting about $T=0.5$, string-like particle exchange motion or ”string-like” motion emerges. In particular, this form of collective motion involving a permutational exchange motion starts becomes conspicuous. Previous simulations on both crystalline materials and glass-forming liquids exhibit what appears to be a superficially similar collective motion so this type of motion to be rather general. Importantly, the $T$ regimes associated with these distinct regimes of particle diffusion and relaxation align with the kink observed in Fig. 10(c) and Fig. 12(c).
We next investigate the particle displacement process in greater detail. In Figure 14, we present an illustration of a putative phason flip at the very low temperature, $T=0.3$. Notably, all observed jumps occur at structural imperfections within the quasicrystal. In the case of a hexagonal tile, nonadjacent particles exist in a relatively unstable state, as predicted by the LJG potential. Consequently, particles apparently tend to move in search of more energetically favored positions, providing an energetic impetus for the phason flips. Figures 14(a) and 14(b) depict the spatial distribution of the same particle cluster at $t=0$ and $t=3\times 10^{4}$, respectively. The red atom undergoes a phason flip (or several reversible phason flips) during this period. Interestingly, regardless of how many times these flips occur, the form of defects, i.e., the hexagonal tiles, persists and no topological transition is observed. Moreover, the number of defects exhibits a strong correlation with the preparation process of the quasicrystal phase, specifically the cooling rate. Quasicrystals with lower potential energy, corresponding to a lower concentration of defects, can be obtained through a slower cooling rate. In practice, forming ”perfect” quasicrystals without defects is challenging due to the limitations of the timescale and length scale of our atomistic simulations, making it difficult to reach an equilibrium concentration of defects.
Figure 14: Illustration of the phason flip mechanism at low temperature $T=0.3$. (a) and (b) are the spatial distribution of a particle cluster at $t=0$ and $t=3\times 10^{4}$, respectively. The yellow arrow represents the displacement vector of the red particle which has been scaled 1.5 times for clear observation.
[…]
Figure 15: Typical string-like motion mechanisms at high temperature $T=0.8$. (a) A single loop. (b) A large loop containing several hexagonal units. (c) A long straight line. (d) A serpentine line. The yellow arrows represent displacement vectors.
Figure 16: Minimum energy pathway of a phason flip. The yellow arrows represent displacement vectors which are scaled twice for clear observation.
Figure 17: Minimum energy pathway of a string-like motion in form of a single loop. The yellow arrows represent the real displacement vectors.
Figure 18: Minimum energy pathway of a string-like motion in form of a ruddged line. The yellow arrows represent the real displacement vectors.
III.2.5 Energy landscape of atomic motion
Given the unique atomic motion mechanisms observed in quasicrystals, we further analyze these two particle displacement modes from an energetic perspective to obtain a deeper insight into this striking phenomenon. We next compute both the activation energy and the corresponding minimum energy pathway (MEP) for both phason flip and string-like motion using the nudged elastic band (NEB) method [115].
First, Figure 16 illustrates the MEP of what we temtatively identify as a phason flip. The displacement process apparently involves an atom climbing over a single energy saddle point. We select specific characteristic points along the MEP to observe the detailed atomic displacement form. As the energy increases, the atom gradually moves into the position of an existing defect. Upon reaching the saddle point F, a new form of defect appears. Subsequently, the atom gradually settles into a new equilibrium position. The activation energy required for this entire process is approximately 2.5, aligning well with the data obtained from the Arrhenius law in Fig. 12 at the low-temperature regime.
In the case of string-like motion, depicted in Figs. 17 and 18, the energetic landscape traversed evidently involves a much greater complexity, the MEP involves multiple saddle points and several meta-stable positions. The illustrative ring-like exchange motion initiates with the movements of two or three particles of the hexagonal ring, accompanied by the generation of a defect left in the tail (Fig. 17D). This forms the first meta-stable state, with $\Delta E$ for this process being about 6, nearly three times the value for a phason flip. Subsequently, it evolves into a semi-loop form (Fig. 17E) with the disappearance of the produced defect in Fig. 17C. This cluster tends to be also meta-stable, causing the MEP to drop to a point F. At this point, the structure and bonds differ slightly from the initial configuration after traveling half the distance of the nearest neighbor separation, with the energy of state F being higher than that in the initial state. No further meta-stable state is encountered after this point in the collective motion evolution, apparently prompting the cluster of atoms to continue its motion in a collective way. Ultimately, a full loop is completed, and the structure returns to its initial equilibrium position.
For an open string, as illustrated in Fig. 18, a similar phenomenon to the ring-like string is observed. This type of string starts with the motion of only two or three particles (B and C), accompanied by the generation and annihilation of defects in arbitrary locations in the string. A near linear particle displacement path is observed at F where every particle moves half the distance of the nearest neighbor separation, culminating in the completion of a full string. At the end of the string ”lifetime”, every atom has traveled the nearest-neighbor distance, and a defect is left as ”debri” from this form of collective motion. Throughout the process, the saddle point appears at point J, with an energy barrier of more than 10.
III.2.6 Dynamics of string-like cooperative motion
In recent years, extensive atomistic simulations have unveiled mechanisms of string-like cooperative motion in amorphous solids, particularly in metallic glasses (MGs), occurring at elevated temperatures around the glass transition or above [95, 116, 117, 118]. This phenomenon has emerged as a characteristic feature of glass-forming liquids generally. In our study, we extend this type of investigations to quantify such motions in the diffusion of the 2D dodecagonal quasicrystal and the relation between these motions to the $T$ dependent activation energy. To initiate our exploration, we provide a concise overview of the methodology employed to quantify such string-like cooperative motion.
The initial step involves identifying the “mobile” particles that accommodate dynamics in the system. Here, mobile particles are defined as those exhibiting a displacement magnitude $r(t)$ greater than the typical amplitude of particle vibration over a specified time interval $t$. Observations of the diffusion mechanism indicate that particles engaged in the string-like motion typically undergo a multistage process to complete a collective motion event so that assigning a definite activation from an analysis of individual stages in this process is difficult. However, the largest changes in $\Delta E$ occur at lower $T$, a trend that parallels the growth of the activation energy upon cooling apparent in the insert to Fig. 10(c).
Next, we employ a mathematical criterion to identify these mobile particles based on a threshold of atomic displacement, specifically we apply the commonly invoked condition $\left|\mathbf{r}_{i}(\Delta t)-\mathbf{r}_{i}(0)\right|>0.6$. In this case, the selected particles have climbed over the energy barrier. Importantly, the count of mobile particles is inherently temperature-dependent, decreasing as the system cools. To identify correlated atomic motion, we employ an established procedure outlined in Ref. [119]. Collective motion implies that the spatial relationship between two particles is retained to some extent as they traverse the system. Specifically, for a pair of “mobile” particles, denoted as particle $i$ and particle $j$, we consider them to be engaged in string-like motion if they persist within each other proximity over time. This proximity relation is defined by the condition $\left|\mathbf{r}_{i}(\Delta t)-\mathbf{r}_{j}(0)\right|<\delta$, where $\delta$ is a parameter chosen to be at least smaller than half of the bond length. In our case, we set $\delta=0.2$. Importantly, we confirm that the results of string-like motion are not significantly influenced by the choice of $\delta$.
Utilizing the methodologies outlined earlier, we extract several crucial parameters, including the number of mobile particles participating in string-like motion and the fraction of strings. In Fig. 19(a), we illustrate the average string length (number of particles) at various temperatures. In alignment with previous observations [119, 120], the average length $L(t)$ is negligible within a very short time (ballistic motion) and progressively increases with time intervals after long-range diffusion occurs, signifying weak correlation of mobile strings. Notably, $L(t)$ exhibits a peak at a characteristic time denoted as $t_{L}$. This timescale defines the characteristic string “lifetime” [97]. The peak value intensifies with cooling, indicating enhanced cooperative mobility at lower temperatures. Simultaneously, the position of the peak shifts to longer times as the temperature decreases, suggesting a deceleration in cooperative string-like diffusion.
Figure 19: Analysis of the string-like cooperative motions. (a) Average string length $L(t)$ as a function of time at different temperatures. The inset shows the $T$-dependence of the peak value. (b) Diffusive relaxation time $t_{L}$ vs $1/T$. The inset shows a linear scaling between $t_{L}$ and $t^{*}$. (c) The distribution of string lengths $P(L)$ at different temperatures, fitted by an exponential law. The inset exhibits the average length of the strings versus temperature.
Upon further examination, it becomes evident that the string lifetime is approximately proportional to $t^{*}$ (inset of Fig. 19(b)), where the latter represents the peak time of the non-Gaussian parameter (Fig. 11(a)). This correlation is consistent with previous observations in other systems [103, 121, 95, 119]. Earlier studies have already established a connection between $t^{*}$ and $D$, i.e., $D/T\sim t^{*}$. The linear scaling relationship can be explained by the decoupling phenomenon. Specifically, it is widely observed that the structural relaxation time experiences a more rapid increase with decreasing temperature compared to the time scale associated with the diffusion coefficient $D$. This leads to the breakdown of the Stokes-Einstein relation in glass-forming liquids between $D/T\sim\tau_{\alpha}$. We will provide further support for this decoupling relation between these quantities in the subsequent analysis of the same form as observed in MG forming materials.
Additionally, the characteristic time $t_{L}$ is slightly longer than the characteristic time $t^{*}$, as illustrated in the inset of Fig. 19(b). Moreover, the time scale of $t_{L}$ is inherently shorter than $\tau_{\alpha}$ due to the breakdown of the Stokes-Einstein relation [103]. The diffusive relaxation times $t_{L}$ and $t^{*}$, occupy an intermediate position between the fast $\beta$ relaxation time and the slow $\alpha$ structural relaxation time.
Finally, to quantify the geometrical properties of the strings, we analyzed the distribution of string length, denoted as $P(L)$, at different temperatures at the characteristic time scale $t_{L}$. The accurate length calculations involved a meticulous three-step process: (1) Filtering particles engaged in string-like motions. (2) Identifying the onset particle of each string from the string particles, where the onset particle does not overlap with the positions of other string particles over the time interval $\Delta t$ of observation. (3) Sequentially selecting string particles starting from the onset particle, allowing us to determine not only the length of each string but also its form (whether it is a closed or open string).
As depicted in Fig. 19(c), the distribution functions $P(L)$ at different temperatures exhibit an approximately exponential decay, as noted in previous work [97, 103, 119]. The deviations form exponential behavior at different $T$ can be described by a power-law variation with an exponential cut-off,
$P\left(L\right)\sim L^{-\theta}\exp\left(-L/L_{0}\right)$ (17)
where $L_{0}$ is a length parameter proportional to the average string length $\langle L\rangle$. This type of power law times an exponential scaling for the distribution of string length has been validated across various systems, including glassing-forming liquids [97] and lipid membranes [122, 123], where $\theta$ exhibits a similar value. Here, we find the exponent $\theta$ increases on cooling, which may reflect a progressive topological transition between open strings at high temperature to gradually increasing closed clusters at low temperature, as observed in our system and previous studies [97, 85, 124].
The inset of Fig.19(c) depicts the temperature dependence of the average string length $\langle L\rangle$. Clearly, the size or average length of the strings increases as $T$ decreases, qualitative accord with the $T$ dependence of $\Delta G$ in Fig. 10(c). We suggest the noise in our $L(T)$ data reflects the uncertainty created by the $T$ dependence of the $\theta$ exponent in Eq. (17). Given equation for $P(L)$ so that a greater effort needs to be made more quantitative estimates of $L(t)$. Our estimates of the $T$ dependence of the activation energy are also a bit noisy and the estimation of this $T$ dependence needs further attention. Qualitatively, however, the growth of the activation free energy upon cooling is accompanied by the scale of collective motion, a signature feature of glass-forming liquids.
III.2.7 Interrelation between $t^{*}$, $\tau_{\alpha}$ and $D$
The growth in the peak height of $\alpha_{2}(t)$ (Fig. 11) reflects the strongly dynamic heterogeneity in the quasicrystal on cooling. As in glass-forming liquids, this quantity exhibits a peak at characteristic time $t^{*}$. The peak value of $\alpha_{2}\left(t^{*}\right)$ increases as the system progressively cooled. As noted before studies have established a robust correlation between $t^{*}$ and $D$ in both atomic glass-forming liquids and heated crystalline [100, 125, 91, 103, 85, 126]. In the following discussion, we also explore whether power-law scaling relation between $D/T$ and the structural relaxation time $\tau_{\alpha}$ or a “decoupling relation” exists in our quasicrystal material.
Figure 20: Scaling of the temperature normalized self-diffusion coefficient $D$ with (a) the non-Gaussian characteristic time $t^{*}$ and (b) the $\alpha$ relaxation time $\tau_{\alpha}$.
We next consider the temperature dependence of the self-diffusion coefficients obtained from the long time limit of the MSD, with $t^{*}$ derived from the variation of the non-Gaussian parameter in $\alpha_{2}(t)$ (see Fig. 20(a)). This comparison reveals a robust power-law relationship between $D/T$ and $t^{*}$, expressed numerically as $D/T\sim(1/t^{*})^{1.03}$. Notably, the scaling exponent is equal to unity to within numerical uncertainty. This scaling evidently resembles the dynamics of metallic glass-forming liquids modeled by embedded atom method (EAM) potentials [102, 85, 103] and the Kob-Anderson model in which a simple binary Lennard-Jones fluid is often studied as ”generic” glass-forming liquid [102]. Notably, this inverse scaling relation was found to not hold in simulations of superionic crystalline ${\rm UO}_{2}$ [100], pointing again to a commonality of the dynamics of our quasi-crystalline system with the dynamics of glass-forming liquids. Moreover, finding establishes the new notion that $t^{*}$ can serve as a distinctive characteristic timescale for diffusion in the quasicrystalline system.
We next explore the quantitative relationship between the characteristic timescale of diffusion $t^{*}$ just established and structural relaxation $\tau_{\alpha}$. Many previous studies of glass-forming liquids have indicated that $t^{*}$ is related to $\tau_{\alpha}$ by a apparently universal scaling relation [91]: $t^{*}\sim({\tau_{\alpha}})^{1-\zeta}$. Here, $\zeta$ defines the “decoupling exponent”, quantifying breakdown of the Stokes-Einstein relationship in glass-forming liquids given that $D/T$ scales as $1/t^{*}$. The case in which $\zeta=0$ corresponds to a situation in which the classical Stokes-Einstein relation applies. This “decoupling” phenomenon is a signature feature of metallic glass-forming liquids [91, 125] and glass-forming liquids more broadly [127], but no decoupling is observed in crystalline materials [85, 100, 126]. This provides a sharply defined criterion for determining whether the relaxation dynamics of quasi-crystalline materials is more like crystalline or glass-forming materials.
Finally, we consider the “fractional Stokes-Einstein relation” in the dodecagonal quasicrystal system. Figure 20(b) shows this scaling relationship holds between $D/T$ and $\tau_{\alpha}$, where we see the specific scaling relation, $D/T\sim(1/\tau_{\alpha})^{0.67}$. The decoupling exponent $\zeta$, for our quasicrystal material, is approximately 0.33 which is rather typical value in glass-forming liquids [127, 91, 125]. More generally, $\zeta$ is found to lie in the range 0.16–0.46. Specifically, previous studies have estimated $\zeta$ as 0.26 in the case of a Pd-Cu metallic glass-forming alloy [128], and 0.18 and 0.17 was reported for the Kob-Anderson binary LJ model in three dimensions [129, 130]. A breakdown of the Stokes-Einstein relation or the decoupling phenomenon is clearly observed in our quasi-crystalline system, a characteristic of glass-forming liquids generally. In contrast, simulations of heated crystalline systems [85, 100, 126] have consistently shown the lack of decoupling to within numerical uncertainty, even these systems show a peak in the non-Gaussian parameter, strings and other features of the dynamics of glass-forming liquids. We have finally been able to find definitive evidence indicating that quasi-crystalline materials being more like metallic glass-forming systems than crystalline materials.
IV Concluding remarks
In summary, our investigation has comprehensively explored the thermodynamic and dynamic attributes of one-component dodecagonal quasicrystals in two dimensions, employing the isotropic LJG pair potential as a benchmark model system. The thermodynamic stability of the dodecagonal quasicrystal is attained under well-defined parameter conditions. Notably, the transition from the liquid phase to the quasicrystal phase is marked by two distinct sharp thermodynamic transitions, although the actual order of these transitions remains to be determined. The consistent alignment of our model system with previous experimental findings supports the generality of our conclusions.
Our examination of the thermodynamic properties and structure entails a comprehensive analysis of density, heat capacity, specific heat, radial distribution function, and static structure factor within the prepared quasicrystal. The quasicrystal exhibits structural characteristics reminiscent of crystalline solids. The RDF of the quasicrystal unveils both short-range order and, to some extent, questionable long-range order, considering the limitations of the LJG pair potential in capturing long-range and many-body interactions.
The dynamic properties of quasicrystals closely parallel those of metallic glass-forming materials, a relationship that is most conspicuous in the dynamical properties of these materials, while their structures appears to have a closer relation to crystals in relation to thermodynamic properties relating to the existence of an ordered state having a translational and rotational symmetries and exhibiting thermodynamic transitions between phases of different types of order. Quasi-crystals really are a hybrid form of matter that challenges our conceptions about the nature between structure and dynamics. The temperature dependence of the structural relaxation time $\tau_{\alpha}$ and $D$ both exhibit a transition in the quasi-crystal regime characterized by kink in Arrhenius curves for $\tau_{\alpha}$ and the diffusion coefficient having a singular upward curvature, a phenomenon observed in rather generally phenomenologically in materials that Angell has classified as intermediate forms of matter [10].
As the quasi-crystal material is cooled, we observe an increase in dynamic heterogeneity paralleling the slowing down of the dynamics. The characteristic time of diffusion coefficient, $t^{*}$, at which the non-Gaussian parameter exhibits a maximum, reflecting the magnitude of the dynamic heterogeneity in the system. Notably, the characteristic time $t^{*}$ is smaller than the structural relaxation time $\tau_{\alpha}$. Our comprehensive analysis of the self-diffusion of quasicrystals takes into account minimum energy pathways, revealing the intricate pathways and intermediate features of string-like motions within the dodecagonal quasicrystals. Although heated crystalline materials and glass-forming liquids exhibit many similar features in their dynamical properties, in the final analysis the dynamic properties of quasicrystals show a greater resemblance to metallic glass-forming materials than heated crystals. Previous neutron scattering studies [12] have repeatedly arrived at the same conclusion based on notably different dynamical properties [10]. Measurements on $D$ in real quasi-crystal materials have also indicated a kink in their Arrhenius curves for the diffusion coefficient [50] as we observe in our simulations of diffusion in our quasi-crystalline material. Our simulations are thus broadly consistent with measurement studies on quasi-crystalline materials.
Another striking aspect of the dynamics of our quasi-crystalline material, shared with glass-forming materials broadly, is that the activation free energy and the string length both increase upon cooling. Our data for both of these quantities exhibit some scattering and a greater effort needs to be made to estimates these properties with less uncertainty. Nonetheless, there is a clear evidence that activation barrier for particle displacement to a distances required for diffusion and structural relaxation clearly grows upon cooling in both our quasi-crystal materials and glass-forming liquids generally. This is apparently a new finding that we did not ourselves initially expect.
It is evident that cooperative string-like collective motion play a pivotal role in the dynamics of quasi-crystalline materials and we expect this to be the case in this class of materials broadly. Over an extended observation period, the string length at different temperatures follows an length distribution similar to previous observations on both glass-forming liquids and crystalline materials. The characteristic time which the string length peaks, $t_{L}$, exhibits a strong correlation with diffusive relaxation time. Additionally, the present quasicrystal exhibits a breakdown of the Stokes-Einstein relationship between $D$ and $\tau_{\alpha}$, akin to many previous observations on glass-forming materials, but not to crystalline materials where this decoupling phenomenon is not observed. Of course, it is not clear why the Stokes-Einstein relation should apply to quasi-crystalline or crystalline materials, which are solids rather than liquids.
Further research is imperative to enhance our understanding of the phason flip motions that we observe and how these motions fit in with the conspicuous string-like collective dynamics, and its interrelation with vibrational properties. In particular, the present understanding of the vibrational properties of quasicrystals is incomplete in our view. For example, the potential existence of an excess vibrational mode, akin to the boson peak observed in amorphous solids and heated crystals, poses intriguing questions. The speculation that phason flips might involve such excess modes merits thorough computational investigation. Unraveling the intricacies of these vibrational phenomena in quasicrystals will undoubtedly contribute to a broader understanding of the dynamics of quasi-crystalline materials. However, we think the pursuit of these questions should be based on a quasi-crystal model in three dimensions.
Acknowledgements.
This paper was partially supported by the National Natural Science Foundation of China (Grant No. 12072344) and the Youth Innovation Promotion Association of the Chinese Academy of Sciences. The numerical calculations in this paper were carried out on the ORISE Supercomputer. M.B. acknowledges the support of the Shanghai Municipal Science and Technology Major Project (Grant No.2019SHZDZX01) and the sponsorship from the Yangyang Development Fund. W.-S.X. acknowledges the support from the National Natural Science Foundation of China (Grant Nos. 22222307 and 21973089).